# Spinning LQG black hole as a particle accelerator

### v1

@@ V Discussions
The observation that center-of-mass energy diverges during particle collision is particularly intriguing from an astrophysical perspective. Supermassive black holes, such as those located at the center of the giant elliptical galaxy M87 and at the center of our galaxy, Sgr A, might accelerate particles through this mechanism, potentially linking it to the observed ultra-high-energy cosmic rays, which can reach up to $10^{20}$ eV. These collisions could also produce exotic massive particles, offering an avenue for further investigation in a realistic setting as an extension of our study. Consequently, we believe that our research could be connected to observational phenomena aimed at probing Planck-scale physics. A potential direction for future work involves calculating the efficiency of energy extraction in the collisional Penrose process for spinning LQG black holes.

### v2

@@ V Discussions
The observation that center-of-mass energy diverges during particle collision is particularly intriguing from an astrophysical perspective. Supermassive black holes, such as those located at the center of the giant elliptical galaxy M87 and at the center of our galaxy, Sgr A, might accelerate particles through this mechanism, potentially linking it to the observed ultra-high-energy cosmic rays, which can reach up to $10^{20}$ eV. These collisions could also produce exotic massive particles, offering an avenue for further investigation in a realistic setting as an extension of our study. Consequently, we believe that our research could be connected to observational phenomena aimed at probing Planck-scale physics. A potential direction for future work involves calculating the efficiency of energy extraction in the collisional Penrose process for spinning LQG black holes. From table 3, we observe an enhancement in center-of-mass energy, which could result in increased efficiency compared to the Kerr black hole.
VI Acknowledgment
ANK’s research was supported by the Croatian Science Foundation Project No. IP-2020-02-9614 Search for Quantum spacetime in Black Hole QNM spectrum and Gamma Ray Bursts.